In a second order reaction 2A is converted into products, the final concentration is .28M. What is the initial concentration if k=.32M-1 s -1 and the time is 5 seconds.


1
Expert's answer
2020-10-14T14:03:32-0400

k=0.32M-1 s -1

time ,t = 5 sec

final concentration, A= 0.28M


 2A--> products 


for 2nd order , integrated rate law is given


1/A0 = 1/A -Kt 


plug in above numbers


1/A0 = 1/0.28 M - 0.32M-1 s -1 * 5 s

  1/A0 = 1.971428 M-1


A0 = 0.507 M


ANSWER

Final concentration = 0.507 M